13) Ann proposes to to invest $150,000 in a business for a 30% shareholding. What is the value of the business?

a. Before the investment: _______

b. After the investment: _________
Answer #1

Suppose the value of the business before investment be $A
Now, 30% of A is 150000 or A*30%=150000
=>A=150000/30%=500000

After the investment, the value will become value before investment + Investment amount=500000+150000=650000

a. Value before the investment= $500000

b. Value after the investment= $650000